OpenStudy (anonymous):

Solve: |x| = 5 Answer {5} {-5} {-5, 5} {-1, 5}

OpenStudy (mathstudent55):

To solve an absolute value equation do the following: 1. Take what's inside the absolute value symbol and set it equal to the right side. 2. Write the word "or". 3. Take the negative of what's inside the absolute value symbol and set it equal to the right side. 4. Solve both equations with "or" in between.

OpenStudy (mathstudent55):

Example: Solve |x - 2| = 8 x - 2 = 8 or -(x - 2) = 8 x = 10 or -x + 2 = 8 x = 10 or -x = 6 x = 10 or x = -6 final answer

OpenStudy (grant330sims):

The answer is A.

OpenStudy (anonymous):

x=5,-5